There are 20 students in a class, 10 who are girls, and 10 who are boys. The teacher is choosing 4 of these students to be on a committee. What is the probability that there will be at least one girl and at least one boy? Use complementary counting.
Answer by stanbon(75887)

There are 20 students in a class, 10 who are girls, and 10 who are boys. The teacher is choosing 4 of these students to be on a committee. What is the probability that there will be at least one girl and at least one boy? Use complementary counting.
-------------------------
Number of ways to fail:
Choose 4 boys: 1 way
Choose 4 girls: 1 way
-------------
# of possible groups of 4: 20C4 = 4045
------
P(at lest one boy and at least one girl) = 1 - 2/4045 = 4043/4035
======================================================
Cheers,
Stan H.
================================